You are offered a court settlement in the following terms: you will receive 9 equal payments of $7928 each every year, with the first payment being made 4 years from now. The current annual interest rate is 4%. Assume yearly compounding. What is this settlement worth in present value terms?
Answer #1

Value at year 3 = Annuity * [1 - 1 / (1 + r)n] / r

Value at year 3 = 7928 * [1 - 1 / (1 + 0.04)9] / 0.04

Value at year 3 = 7928 * [1 - 0.70259] / 0.04

Value at year 3 = 7928 * 7.43533

Value at year 3 = 58,947.309

Present value = FV / (1 + r)n

Present value = 58,947.309 / (1 + 0.04)3

Present value = 58,947.309 / 1.124864

Present value = $52,403.94